OpenStudy (keana):

What is the value of the coefficient b when the quadratic equation y=(3x-5)(2x+3) is written in standard form? A: 19 B: 6 C: -15 D: -1

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Pac1f1cIslander You are wrong unfortunately.

OpenStudy (anonymous):

\[(3x - 5)(2x + 3) = 3x(2x + 3) - 5(2x + 3) = 6x^{2} + 9x - 10x - 15 = 6x^{2} - x - 15\]This is in the form:\[ax^{2} + bx + c\]Which is the b-term?

OpenStudy (keana):

Um...

OpenStudy (anonymous):

What's cut off should be \[6x^{2} - x - 15\]

OpenStudy (keana):

I really don't know ...

OpenStudy (anonymous):

bx = -x What is b?

OpenStudy (keana):

-1 ?

OpenStudy (anonymous):

No keana, 6 Option B is 6

OpenStudy (anonymous):

It's definitely -1.